OpenStudy (anonymous):

On the coordinate grid of a map. Josie's house is located at (2,7). Her school is located at (-5,5). If each map unit equals one mile, what is the approximate distance from her house to school? A) 2.83 miles B) 4.79 miles C) 7.28 miles D) 14.76 miles

OpenStudy (anonymous):

Let A(2,7) and B(-5,5) then dist AB = Sqrt((2+5)^2+(7-5)^2)= sqrt (53)= 7.28 as each unit is 1 mile so the distance is 7.28 miles